OpenStudy (anonymous):

Find the equation of the ellipse with center (2,0), focus (5,0), b=4..

OpenStudy (tkhunny):

You are given b and c. What is the relationship to find a?

OpenStudy (anonymous):

you know it is going to look like \[\frac{(x-2)^2}{a^2}+\frac{y^2}{b^2}=1\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you are told \(b=4\) so it must be \[\frac{(x-2)^2}{a^2}+\frac{y^2}{16}=1\]

OpenStudy (anonymous):

yes yes..

OpenStudy (anonymous):

where's focus? hehe

OpenStudy (anonymous):

the focus evidently is \((5,0)\) because you are told that is one of them since it is 3 units to the right of \((2,0)\) the other is 3 units to the left, i.e. at \((-1,0)\)

OpenStudy (anonymous):

this tells you that \(c=3\) and so \(a^2=b^2+c^2\) gives \(a^2=16+9=25\)

OpenStudy (anonymous):

final answer is therefore \[\frac{(x-2)^2}{25}+\frac{y^2}{16}=1\]

OpenStudy (anonymous):

shall we check it?

OpenStudy (anonymous):

i guess the answer is 16x^2 + 25y^2 - 64x -336??

OpenStudy (anonymous):

i like my answer

OpenStudy (tkhunny):

@satellite73 's answer is particularly compelling since the last post of @silverxx is incorrect. I didn't check the arithmetic, but it isn't an equation. That's just not right.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

aaaah okay thank you so much @tkhunny @satellite73 :))